OpenStudy (anonymous):

Which is more likely, to throw at least 1 six with 6 dice, or at least 2 sixes with 12 dice, or at least 3 sixes with 18 dice?

OpenStudy (anonymous):

Try it out and see what happens.

OpenStudy (anonymous):

For the first one it is: \[ \binom{6}{1}\left(\frac{1}{6}\right)^1\left(\frac{5}{6}\right)^5 \]The second one is: \[ \binom{6}{2}\left(\frac{1}{6}\right)^2\left(\frac{5}{6}\right)^{10} \]The third one is: \[ \binom{6}{3}\left(\frac{1}{6}\right)^3\left(\frac{5}{6}\right)^{16} \]

OpenStudy (anonymous):

I don't get it. How did you get this?

OpenStudy (anonymous):

And I think you are wrong, first one equals about .4, second one equals about .067, last one equals about .006 . this is what I got on calculator. It doesn't make sense.

OpenStudy (anonymous):

Oooh wait

OpenStudy (anonymous):

Should be \[ \binom{12}{2} \]for the second and \[ \binom{18}{3} \]for the third...

OpenStudy (anonymous):

Basically, the fraction parts are finding out the probability that the 6's come up consecutively at the start and then the rest are not 6's.

OpenStudy (anonymous):

Then the binomials count how many ways you could shuffle the 6's around, so you account for every sequence they could appear in.

OpenStudy (anonymous):

Hmmm, it's looking a lot like the binomial distribution but...

OpenStudy (anonymous):

But it seems like it actually is the binomial distribution.